The solubility of michael amine is 25 ઈ C. The temperature is 8.49 times higher in water than in chloroform. Determine the percentage of base remaining in the following cases in 1000 ml, chloroform. (a) Four times, 200 ml each time. Extraction from water. (b) Extraction from 400 ml of water twice, each time. (c) Extraction from 800 ml of water once.


1
Expert's answer
2021-04-05T07:45:00-0400

"f_n=(\dfrac{V_1}{DV_2+V_1})^n"

"D = 8.49"

If we have 1 L of water and 800 ml of chloroform:

a) "f_4=(\dfrac{1000}{8.49\times200+1000})^4=0.01887 = 1.887" %

b)"f_2=(\dfrac{1000}{8.49\times400+1000})^2=0.05175 = 5.175" %

c)"f_1=(\dfrac{1000}{8.49\times800+1000})=0.1283=12.83" %